OpenStudy (anonymous):

if a = 2, find the value of: a+3

OpenStudy (anonymous):

5

OpenStudy (aroub):

substitute 2 by a 2+3=5

OpenStudy (chaise):

if a = 2 a+3 = ? 2+3 = 5

OpenStudy (anonymous):

No, no, no, please, no more....

OpenStudy (anonymous):

i am going to forget vectors cones progressions integers derivatives with this